Calculate the momentum of a photon of light of wavelength 500 nm.

Grade:11

1 Answers

Deepak Patra
9 years ago
Sol. λ = h/p ⇒ P = h/λ = 6.63 *10^-34/500 * 10^-9 J-S = 1.326 * 10^-27 = 1.33 * 10^-27 kg – m/s.